OpenStudy (anonymous):

Will Fan And Medal Bailey was building a pond. The pond held 40.5 gallons of water. She had already put 27.9 gallons of water in it. Bailey said she only needed to add 2.6 gallons to fill it. Is this a reasonable estimate?

OpenStudy (anonymous):

A. Yes, it is reasonable because 27.9 is about 28, and 40.5 is about 41. 41 – 28 = 3 so 2.6 gallons is a reasonable estimate. B. No, it is not reasonable because 27.9 is about 28, and 40.5 is about 41. 41 – 28 = 13, so 13 gallons is a more reasonable estimate. C. No, it is not reasonable because 27.9 is about 30, and 40.5 is about 50. 50 – 30 = 20, so 20 gallons is a more reasonable estimate.

OpenStudy (phi):

any thoughts on this one ?

OpenStudy (anonymous):

Well at first I wanted to say it was C.

OpenStudy (phi):

They are rounding the numbers to the nearest integer. Do you know how to round 40.5 to the nearest integer ?

OpenStudy (anonymous):

not correctly

OpenStudy (phi):

the integer position is the "ones" position when we look at 40.5 we need to decide is it closer to 40 or to 41 the rule is: look at the next digit to the right (in this case the tenths digit). if 5 or more, round up, else just drop the decimals

OpenStudy (anonymous):

5 is rounded up to 41

OpenStudy (phi):

and 27.9 is rounded to either 27 or 28 ?

OpenStudy (anonymous):

28

OpenStudy (phi):

so (roughly) the pools has 28 gallons in it, and we need 41 to fill it up. how many more gallons do we need?

OpenStudy (anonymous):

41 - 28 = 13.

OpenStudy (phi):

is that enough to answer the question?

OpenStudy (anonymous):

yes, thanks.